OpenStudy (jaclynsanc557):

What is the solution to the equation -(5-(a+1))=9-(5-(a-3))? a=-5 a=-3 a=5 a=3

OpenStudy (phi):

it looks like there is a typo in this question, because there is no solution

OpenStudy (jaclynsanc557):

-(5-(a+1))=9-(5-(2a-3)) ** sorry

OpenStudy (jaclynsanc557):

@phi

OpenStudy (phi):

in that case, I would use the distributive law to "expand" the expression. Start with -(a+1) which means -1 * (a+1) multiply -1 times each term inside the parens. what do you get ?

OpenStudy (jaclynsanc557):

so i would multiply a=1 by -1 and 2a-3 ny a -1 also?

OpenStudy (jaclynsanc557):

a+1 *

OpenStudy (phi):

we don't know what "a" is but to show multiplying by -1 we would write -1 * a (that means -1 times a) we could simplify that to just -a (which is how people usually would write it)

OpenStudy (phi):

what do you get for -1 * (a+1) ?

OpenStudy (jaclynsanc557):

a -1? i'm so bad at this. the numbers get tangled up in my head and i get confused.

OpenStudy (phi):

Here is an example, using letters: a(b+c) which is a*(b+c) the distribute law says: you can multiply "a" times each thing inside in this case: a*b + a*c or in algebra, we can leave off the multiply sign, and write ab + ac so we change a(b+c) into ab+bc try to do the same thing for this -1(a+1) the rule is: write -1 times each term inside the parens, can you try ?

OpenStudy (jaclynsanc557):

-(5-1(a+1))=9-(5-1(a-3)) so then it would look like this? or is that still wrong? sorry i can't get it :(

OpenStudy (phi):

we are doing the problem -(5-(a+1))=9-(5-(2a-3)) and we start by looking at only this small part: -(a+1) (which is part of the left-hand side: -(5-(a+1)) ) and -(a+1) means -1 * (a+1) you should try to learn how to write down: -1 times a and -1 times 1

OpenStudy (jaclynsanc557):

the answer is -5

OpenStudy (phi):

the answer is knowing how to show a= -5

OpenStudy (jaclynsanc557):

yes but i dont get it! im trying to follow you and do the steps but its not clicking.